Please Help! a)Graph 4x+3y=12 b)Give the intercepts c)Give the slope
Answer by checkley71(8403)

3y=-4x+12 or y=-4/3x+4 or thus slope=-4/3 and the y intercept is 4 & the x intercept is 0=-4/3x+4 or -4/3x=-4 or x=3
sorry don't have graphing tools.